OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2 + bx + c = 0 Solution 10 , only solution

OpenStudy (anonymous):

\[(x-10)^2=0\]Expanding:\[x^2-20x+100=0\]

OpenStudy (anonymous):

When you have one and only one solution for a quadratic, the discriminant must be equal to zero (the discriminant is the thing under the square root sign in the quadratic formula). Notice this is the case here:\[\sqrt{b^2-4ac}=\sqrt{(-20)^2-4(1)(100)}=\sqrt{400-400}=0\]